OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7.

OpenStudy (anonymous):

so we know that perpendicular lines have opposite reciprocal slopes. the slope of one of the lines is 1/3 the reciprocal would be 3/1=3 and opposite makes it -3 so we know that the slope of the line perpendicular to it will be -3. use the equation: y-y1=m(x-x1) y-(-2)= -3(x-4) y+2=-3x+12 y=-3x+10 So the line perpendicular to it would be y=-3x+10

OpenStudy (anonymous):

thank you so much! @lacrosse